# A Sanity Check for AI-generated Image Detection

### v1

@@ 5.4 Robustness to Unseen Perturbations
In real-world scenarios, images inevitably encounter unseen perturbations during transmission and interaction, which increases the difficulty in detecting AI-generated images. Herein, we evaluate the performance of different methods in handling potential unseen perturbations i.e., JPEG compression (Quality Factor (QF) = 95, 90) and Gaussian blur ($\sigma$ = 1.0, 2.0). As shown in Table 5, all methods are experiencing a decline in performance due to the disruption of the pixel distribution. This disruption diminishes the discriminative artifacts left by the generative model, making it more difficult to discern between real and AI-generated images. Consequently, the ability of these detectors to robustly identify AI-generated images is significantly compromised. From this table, our method still achieves SOTA performance in both perturbation scenarios, achieving 75.54% (QF = 95) and 81.88% ($\sigma=1.0$). Despite the challenging conditions, our method consistently outperforms other approaches, maintaining a relatively higher Acc in detecting AI-generated images. This superior performance can be attributed to our model’s ability to effectively capture and utilize multi-perspective features, i.e., semantics and noise, between real and fake images, even when the pixel distribution is distorted.
Table 6: Ablation study of different modules in our method.
Module Average
@@
5.5 Ablation Studies
Our method focuses on detecting AI-generated images with mixture of experts, namely patchwise feature extraction (PFE-H and PFE-L for high-frequency and low-frequency patches, respectively) and semantic feature extraction (SFE). These modules collectively contribute to comprehensively identifying AI-generated images from different perspectives. Herein, we conduct ablation studies on each module on $\mathcal{B}_{1}$ and draw the following conclusions from Table 6: (i) The first 3 rows show the results of ablating each module individually, indicating that while these features are useful for detection, they still suffer from inferior performance in accuracy (around 75% Acc). This aligns with our analysis that finding a universal fingerprint for detection is challenging, and multiple features should be considered simultaneously. (ii) The second 3 rows sequentially combine two modules, and this integration of different features can improve performance to some extent. Notably, PFE-L + SFE achieves the best result among the 3 combinations. This result aligns with the motivation of recent SOTA methods, UnivFD and PatchCraft, highlighting the importance of low-frequency patches and semantic features in AI-generated image detection. (iii) On top of above findings, our method with mixture of the three experts achieves the SOTA performance, proving that the simultaneous use of all three features can achieve excellent results. This comprehensive extraction captures subtle discrepancies and underscores the effectiveness of our ensemble method in detecting AI-generated images.
@@ 6 Conclusion
In this paper, we have conducted a sanity check on detecting AI-generated images. Specifically, we re-examined the unreasonable assumption in existing training and testing settings and suggested new ones. In terms of benchmarks, we propose a novel, challenging benchmark, termed as Chameleon, which is manually annotated to challenge human perception. We evaluate 9 off-the-shelf models and show that all detectors suffered from huge performance drops. In terms of architecture, we propose a simple yet effective model, that simultaneously incorporates low-level patch statistics and high-level semantics for AI-generated image detection. Despite our approach demonstrates state-of-the-art performance on existing (AIGCDetectBenchmark [62] and GenImage [74]) and our proposed benchmarks (Chameleon), it leaves significant room for future improvement.
Appendix A Experimental Details
A.1 Implementation Details
Our method includes two key modules: Patchwise Feature Extraction (PFE) and Semantic Feature Embedding (SFE). For PFE channel, we first patchify each image into patches and the patch size is set to be $N=32$ pixels. Then these patches are sorted using our DCT Scouring module with $K=6$ different band-pass filters in the frequency domain. Subsequently, we select two highest-frequency and two lowest-frequency patches using the calculated DCT scores. These selected patches are then resized to $256\times 256$ and extracted their noise pattern using SRM [21]. For SFE channel, we use the pre-trained OpenCLIP [31] to extract semantic features. We adopt data augmentations including random JPEG compression (QF $\sim\text{Uniform}(30,100)$) and random Gaussian blur ($\sigma\sim\text{Uniform}(0.1,3.0)$) to improve the robustness of detectors. Each augmentation is conducted with 10% probability. During the training phase, we use AdamW optimizer with the learning rate of $1\times 10^{-4}$ in $\mathcal{B}_{1}$ and $5\times 10^{-4}$ in $\mathcal{B}_{2}$, respectively. The batch size is set to $32$ and the model is trained on 8 NVIDIA A100 GPUs for only 5 epochs. Our method trains very quickly, only 2 hours are sufficient.
A.2 Baseline Detectors
We choose a set of representative methods in AI-generated detection as baselines for comparison, including frequency-based [20, 37, 70], gradient-based [59], semantic-based [49], reconstruction-based [63], etc.
  • CNNSpot (CVPR’2020) [62] proposes that a naïve image classifier with simple data augmentations (i.e., JPEG compression and Gaussian blur) can generalize surprisingly to images generated by unknown GAN-based architectures.
  • FreDect (ICML’2020) [20] observes significant artifacts in the frequency domain of GAN-generated images and makes use of these artifacts for classification.
  • Fusing (ICIP’2022) [32] designs a two-branch model to fuse global spatial information and local informative features for training the classifier.
  • GramNet (CVPR’2020) [40] leverages global image texture representations to improve the robustness and generalization in detecting AI-generated images.
  • LNP (ECCV’2022) [37] proposes to extract the noise pattern of images with a learnable denoising network and uses noise patterns to train a classifier.
  • LGrad (CVPR’2023) [59] employs gradients computed by a pretrained CNN model to present the generalized artifacts for classification.
  • UnivFD (CVPR’2023) [49] uses CLIP features to train a binary liner classifier.
  • DIRE (ICCV’2023) [63] observes obvious differences in discrepancies between images and their reconstruction by DMs and uses this feature to train a classifier.
  • PatchCraft (Arxiv’2024) [70] compares rich-texture and poor-texture patches from images and extracts the inter-pixel correlation discrepancy as a universal fingerprint for classification.
A.3 Statistics of Public Benchmarks
Table 7 provides a detailed explanation of Benchmark 1 & 2 introduced in our main paper. There are two main benchmarks here: AIGCDetectBenchmark [62] and GenImage [74]. AIGCDetectBenchmark: It is trained on ProGAN and then tested on 16 different test sets, including data generated by both GAN and Stable Diffusion models. GenImage: It is trained on Stable Diffusion V1.4 and tested on a large amount of data generated by Stable Diffusion, with only a small portion of GAN data included. The test sets related to Stable Diffusion in AIGCDetectBenchmark are consistent with those used in GenImage.
Table 7: Statistics of Benchmark 1 & 2. SD and WFIR refer to Stable Diffusion and whichfaceisreal, respectively. The term "Number" only counts on fake images and an equal number of real images is added for each generative model from the same source. The BigGAN test sets in $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ are different, from ForenSynths [62] and GenImage [74], respectively.
Benchmark 1 Benchmark 2
Generator Image Size Number Source Generator Image Size Number Source
Train ProGAN [33] $256\times 256$ 360.0k LSUN [67] SD v1.4 [6] $512\times 512$ 324.0k ImageNet [16]
 Test ProGAN [33] $256\times 256$ 8.0k LSUN [67] BigGAN [13] $256\times 256$ 12.0k ImageNet [16]
StyleGAN [34] $256\times 256$ 12.0k LSUN [67]
BigGAN [13] $256\times 256$ 4.0k ImageNet [16] ADM [17] $256\times 256$ 12.0k ImageNet [16]
CycleGAN [72] $256\times 256$ 2.6k ImageNet [16]
StarGAN [15] $256\times 256$ 4.0k CelebA [39] Glide [46] $256\times 256$ 12.0k ImageNet [16]
GauGAN [50] $256\times 256$ 10.0k COCO [36]
StyleGAN2 [35] $256\times 256$ 15.9k LSUN [67] Midjourney [4] $1024\times 1024$ 12.0k ImageNet [16]
WFIR [9] $1024\times 1024$ 2.0k FFHQ [34]
ADM [17] $256\times 256$ 12.0k ImageNet [16] SD v1.4 [6] $512\times 512$ 12.0k ImageNet [16]
Glide [46] $256\times 256$ 12.0k ImageNet [16]
Midjourney [4] $1024\times 1024$ 12.0k ImageNet [16] SD v1.5 [6] $512\times 512$ 16.0k ImageNet [16]
SD v1.4 [6] $512\times 512$ 12.0k ImageNet [16]
SD v1.5 [6] $512\times 512$ 16.0k ImageNet [16] VQDM [24] $256\times 256$ 12.0k ImageNet [16]
VQDM [24] $256\times 256$ 12.0k ImageNet [16]
Wukong [10] $512\times 512$ 12.0k ImageNet [16] Wukong [10] $512\times 512$ 12.0k ImageNet [16]
DALLE 2 [54] $256\times 256$ 2.0k ImageNet [16]
Table 8: Benchmark1. The average precision (AP %) of different baselines (rows) in detecting real and fake images from different generators (columns). GAN-Average and DM-Average are averaged over the first 8 and the last 8 test sets, respectively. The best result and the second-best result are marked in bold and underline, respectively.
Generator CNNSpot [62] FreDect [20] Fusing [32] GramNet [40] LNP [37] LGrad [59] UnivFD [49] DIRE-G [63] DIRE-D [63] PatchCraft [70] Ours
ProGAN [33] 100.00 99.99 100.00 100.00 99.99 100.00 99.08 58.79 100.00 100.00 100.00
StyleGAN [34] 99.83 88.98 99.50 99.23 98.60 98.31 91.74 56.68 97.56 98.96 99.99
BigGAN [13] 85.99 93.62 90.70 81.79 84.32 92.93 75.25 46.91 99.27 99.42 94.44
CycleGAN [72] 94.94 84.78 95.50 95.33 92.83 95.01 80.56 50.03 99.80 85.26 99.89
StarGAN [15] 99.04 99.49 99.80 99.23 100.00 100.00 99.34 40.64 99.37 100.00 99.99
GauGAN [50] 90.82 82.84 88.30 84.99 78.85 95.43 72.15 47.34 99.98 81.33 97.69
StyleGAN2 [35] 99.48 82.54 99.60 99.11 99.59 97.89 88.29 58.03 97.90 97.74 99.96
WFIR [9] 99.85 55.85 93.30 95.21 91.45 57.99 60.13 59.02 96.73 95.26 99.27
ADM [17] 75.67 61.77 94.10 73.11 94.20 72.95 85.84 99.79 86.81 93.40 98.77
Glide [46] 72.28 52.92 77.50 66.76 88.86 80.42 78.35 99.54 83.81 94.04 98.94
Midjourney [4] 66.24 46.09 70.00 56.82 76.86 71.86 61.86 97.32 74.00 96.48 88.13
SD v1.4 [6] 61.20 37.83 65.40 59.83 94.31 62.37 49.87 98.61 86.14 99.06 98.26
SD v1.5 [6] 61.56 37.76 65.70 60.37 93.92 62.85 49.52 98.83 85.84 99.06 98.20
VQDM [24] 68.83 85.10 75.60 61.13 87.35 77.47 54.57 98.98 96.53 96.26 99.27
Wukong [10] 57.34 39.58 64.60 55.62 92.38 62.48 55.38 98.37 91.07 97.54 98.62
DALLE2 [54] 53.51 38.20 68.12 49.82 96.14 82.55 74.48 99.71 63.04 99.56 99.41
GAN-Average 96.24 86.01 95.84 94.36 93.20 92.20 83.32 52.18 98.83 94.75 98.90
DM-Average 64.58 49.91 72.63 60.43 90.50 71.62 63.73 98.89 83.41 96.93 97.45
Average 80.41 67.96 84.23 77.40 91.85 81.91 73.53 75.54 91.12 95.84 98.18
Appendix B More Experimental Results
B.1 AP Result
We additionally provide classification results regarding AP in Table 8. It is important to highlight that the AP (Average Precision) metric emphasizes different aspects compared to Acc (Accuracy). While Acc focuses on the overall correctness of predictions across all samples, AP provides a more comprehensive evaluation of a model’s performance across various thresholds, particularly in handling imbalanced datasets. On top of that, our method still achieves SOTA performance among these baselines on AP metric, which underscores the superiority of our approach. This indicates that our method not only excels in general prediction accuracy but also maintains robust performance across different decision thresholds, demonstrating its effectiveness in distinguishing between classes even in challenging scenarios.

### v2

@@ 5.3 Robustness to Unseen Perturbations
In real-world scenarios, images often encounter unseen perturbations during transmission and interaction, complicating the detection of AI-generated images. Here, we assess the performance of various methods in handling potential perturbations, such as JJPEG compression (Quality Factor (QF) = 95, 90) and Gaussian blur ($\sigma$ = 1.0, 2.0). As illustrated in Table 6, all methods exhibit a decline in performance due to disruptions in the pixel distribution. These disruptions diminish the discriminative artifacts left by generative models, complicating the differentiation between real and AI-generated images. Consequently, the robustness of these detectors in identifying AI-generated images is significantly compromised. Despite these challenging conditions, our method consistently outperforms others, maintaining a relatively higher accuracy in detecting AI-generated images. This superior performance is attributed to our model’s ability to effectively capture and leverage multi-perspective features, semantics and noise, even when the pixel distribution is distorted.
5.4 Ablation Studies
Figure 4: Visualization of the effectiveness of PFE and SFE Modules. The absence of our semantic feature extraction module results in numerous AI-generated images exhibiting pronounced semantic errors that are incorrectly classified as real. Similarity, the second row demonstrates that when the patchwise feature extraction module is omitted, many AI-generated images, despite lacking semantic errors, contain subtle underlying noise that also leads to their misclassification as true.
Our method focuses on detecting AI-generated images with mixture of experts, namely patchwise feature extraction (PFE-H and PFE-L for high-frequency and low-frequency patches, respectively) and semantic feature extraction (SFE). These modules collectively contribute to comprehensively identifying AI-generated images from different perspectives. Herein, we conduct extensive experiments to investigate the roles of each module on $\mathcal{B}_{1}$.
Table 7: Ablation studies on Patchwise Feature Extraction and Semantic Feature extraction of AIDE.
Module Mean
@@ 5.4 Ablation Studies
Patchwise Feature Extraction. As shown in Table 7, removing either the high-frequency or the low-frequency patches results in obvious performance degradation in terms of accuracy. Without the high-frequency patches, the proposed method is unable to discern that the high-frequency regions of AI-generated images are smoother than those of real images, resulting in performance degradation. Similarly, without the low-frequency patches, the method cannot extract the underlying noise information, which is crucial for identifying AI-generated images with higher fidelity, leading to incorrect predictions.
Semantic Feature Extraction. As shown in Table 7, the performance degrades significantly (76.70% vs 92.77%) when we remove the semantic branch. Intuitively, if the branch for semantic information extraction is absent, our method struggles to effectively capture images with semantic artifacts, resulting in significant drops.
Visualization. To vividly demonstrate the effectiveness of our modules patchwise feature extraction (PFE) and semantic feature extraction (SFE), we conducted a visualization, as depicted in Fig. 4. In first row, the absence of semantic feature extraction results in many images with evident semantic errors going undetected. Similarly, the second row shows that, without patchwise feature extraction, numerous images lacking semantic errors still contain differing underlying information that remains unrecognized. Overall, our method, AIDE, achieves the best performance.
@@ 6 Conclusion
In this paper, we have conducted a sanity check on detecting AI-generated images. Specifically, we re-examined the unreasonable assumption in existing training and testing settings and suggested new ones. In terms of benchmarks, we propose a novel, challenging benchmark, termed as Chameleon, which is manually annotated to challenge human perception. We evaluate 9 off-the-shelf models and demonstrate that all detectors suffered from significant performance declines. In terms of architecture, we propose a simple yet effective detector, AIDE, that simultaneously incorporates low-level patch statistics and high-level semantics for AI-generated image detection. Despite our approach demonstrates state-of-the-art performance on existing (AIGCDetectBenchmark (Zhong et al., 2023) and GenImage (Zhu et al., 2024)) and our proposed benchmark (Chameleon) compared to previous detectors, it leaves significant room for future improvement.
Potential societal impacts. Given that Chameleon demonstrates the capability to surpass the ”Turing Test”, there exists a significant risk of exploitation by malicious entities who may utilize AI-generated imagery to engineer fictitious social media profiles and propagate misinformation. To mitigate it, we will require all users of Chameleon to enter into an End-User License Agreement (EULA). Access to the dataset will be contingent upon a thorough review and subsequent approval of the signed agreement, thereby ensuring compliance with established ethical usage protocols.